OpenStudy (anonymous):

What is the y intercept of a line that passes through (1, 2) and has a slope of - 5/3

OpenStudy (bakonloverk):

\[y-2=\frac{ 5 }{ 3 }(x-1)\]

OpenStudy (bakonloverk):

\[y-2=\frac{ 5 }{ 3 }x-\frac{ 5 }{ 3 }\]

OpenStudy (bakonloverk):

then add 2 to the 5/3 and the -5/3

OpenStudy (anonymous):

thank you

OpenStudy (bakonloverk):

Your welcome